OpenStudy (anonymous):

Solve the equation. |10x – 14| - 5 = 156

OpenStudy (anonymous):

10x - 14 = 160 and 10x -14 = -160

OpenStudy (anonymous):

solve for x. you should get 2 answers

OpenStudy (campbell_st):

well not quite add 5 to both sides of the equation and you get |10x - 14| = 161 then you have 2 equations to solve since its absolute value 10x - 14 = 161 and 10x - 14 = -161 solve eaxh equation for x, for the solutions.